The Ka for benzoic acid is 6.5 x 10-5
.
Calculate the pH of a 0.15 M benzoic acid
solution.
1
Expert's answer
2020-12-01T14:23:55-0500

Ka= 6.5 x 10-5

[C6H5COOH]= 0.15M

C6H5COOH:::::::: C6H5COO- + H+

The equilibrium concentrations are

[C6H5COOH] = 0.15 - x

[C6H5COO-] = x

[H+] = x

Now,

Ka= [C6H5COO-][H+]/[C6H5COOH]

6.5x10-5 = x2/0.15-x

Let's make an approximation 0.15-x is equal to 0.15 and then check for the validity of our approximation

x2/0.15 = 6.5x10-5

x2= 4.33 x 10-4

x= 0.021M

Let's check for the validity of our approximation

Percent ionization of benzoic acid= x/[C6H5COOH] x 100%

= 0.021/0.51 x 100%

= 14%

Our approximation is invalid, approximations of above 5% ionization are generally invalid, therefore we need to set up a quadratic equation.

6.5x10-5 = x2/ 0.15-x

x2= 6.5x10-5 (0.15-x)

x2 = 9.75x10-6-6.5x10-5x

x2 + 6.5x10-5x-9.75x10-6= 0

Using the formula

x= -b+- √b2 - 4ac/2a

Where a= 1, b= 6.5 x10-5, c= 9.75x10-6

x= -6.5x10-5 +- √(6.5x10-5)² - (4)(1)(-9.75x10-6)/2(1)

x= -6.5 x 10-5 +- 6.28x 10-3/2

x= 3.11 x 10-3

Therefore, [H+] = 3.11 x 10-3M

pH= -log[H+]

pH = -log[3.11x10-3]

pH= 2.51